3. Consider a retailer selling blenders currently priced at $54. Suppose it pays $29 per blender from the manufacturer.

A. Initial contribution margin is $54-$29= $25

B. Suppose it is considering a 33% cut in price to boost sales. What is the break-even change in sales required to maintain its profitability?

C. Alternatively, suppose an expert tells the retailer that it should consider raising its price of the blenders to $59 to improve profit. What is the break-even change in sales permissible to again maintain its profitability?

D. Using the break-even change in sales you obtained in (b) and (c), plot the break-even curve for the retailer.

E. Suppose the retailer’s market research team determines that the elasticity of demand for consumers of blenders is – 1.5. What does this imply about the actual demand for blenders in case of the two situations: a 33% price cut or a price increase to $59? Plot the demand curve alongside the break-even curve to show the difference between the two curves.

F. Can you make recommendations to the retailer regarding which strategy makes more sense: a 33% price cut or a price rise to $59 from its current price level of $54?
Answer #1
Assuming Current Sales in units=N
Fixed Costs=F
Current Profit=N*25-F
Break even Sales =N1
Changed Price =(1-0.33)*54= $36
Changed Contribution Margin=36-29= $7
Changed Profit =N1*7-F
For Breakeven,
N1*7-F=N*25-F
7N1=25N
N1=(25/7)*N=3.57N 3.571429
Break even Increase in sales Required 257.14% (3.571429-1)
C Increasing Price to $59
Assume Breakeven Sales=N2
Changed Contribution =59-29=$30
Changed Profit=N2*30-F
N2*30-F=N*25-F
30N2=25N
N2=(25/30)*N=0.83333N 0.833333
Breakeven DECREASE in Sales 0.16667 (1-0.83333)
Breakeven DECREASE in Sales 16.67%